the famous hope diamond at the smithsonian national museum of natural history weighs 45.52 carats. Diamond is a crystallized form of carbon.
a) How many moles of carbon are in the Hope diamond. (1carat = 200mg)
b) how many carbon atoms are in the diamond?
1
Expert's answer
2016-02-16T00:00:57-0500
m(C)=0.2*45,52=9.104g
N(C)=9.104/12.011=0.758 moles
N(C)=0.758*6.02e23=4.56e23 atoms